# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 3.2.3 AI-Based Network Parameter Optimization
In wireless communication systems, parameter configuration profoundly impacts network performance. One of the most important aspects is the appropriate setup of the network parameters, as these determine the design, energy consumption, and performance of the 5G network and need to be aligned with the real-time user distribution and electromagnetic environments. According to the characteristics of wireless networks, network parameters can be divided into two categories: BS and threshold. Specifically, the BS parameters involve switching essential functions, such as VONR, paging, and slicing. The threshold parameters configure thresholds for cell reselect, event decision, and neighborhood switching. Table 6 summarizes various parameters in the current network.
The traditional parameter optimization mainly relies on road test collection and expert experience judgment. The current network parameter optimization methods are as follows: The particular network operator drives a radio monitoring vehicle to gather and report network performance indicators, such as network throughput and audio and video transmission delay, when the network operator determines that there are evident flaws in the communication service quality provided. After that, the operator creates a scatter plot to characterize the network performance further, and lastly, using expert experience, analyzes and optimizes the network parameters [86]. With the network structure’s complexity, traditional parameter optimization’s efficiency is inefficient. The channel is no longer a straightforward one-dimensional circuit loss structure due to the significant increase in 5G antennas; instead, it now extends to the N-dimensional angular power spectrum, producing an extremely large-scale parameter combination. Expert experience-based parameter tuning efficiency is low, making modifying hundreds of millions of 5G network parameters challenging. In addition, traditional empirical tuning methods often rely on inherent rules and cannot fully capture and adapt to complex relationships in the network [87].
@@
However, AI models’ non-linear fitting and large-scale real-time tuning capabilities show unprecedented advantages [84]. First, AI models are better suited for large-scale parameter adjustment. As the number of 5G antennas and network complexity rise, there are an increasing number of possible combinations for network parameters. AI can quickly search and optimize these combinations. Secondly, AI models can better deal with complex nonlinear problems: 5G networks have a highly complex architecture and nonlinear characteristics [88]. By learning and analyzing a large amount of network data, AI technology can better understand the non-linear characteristics of the network to find better parameter configurations. Finally, the AI models align with the real-time requirements of network parameters: in 5G networks, the dynamic changes of user needs and network environment are fast, and network parameters must be adjusted in real time to meet the needs of different scenarios. AI technology can promptly optimize network parameters according to real-time data and scene requirements for better performance.
To optimize network parameters, deep reinforcement learning can be applied [89]. First, each performance indicator of the current communication networks is regarded as a different reinforcement learning state. The model’s network parameter adjustment is viewed as performing various actions. By constantly adjusting the behavior of the reinforcement learning model in other states, the model gets feedback from the environment to know how good or bad each action is. Reinforcement learning models tend to reinforce actions with high rewards. As the rewards overhead is stabilized, deep reinforcement learning gradually converges, thus determining the optimal configuration of the wireless network parameters.

### v3

@@ 2.2.3 AI-Based Network Parameter Optimization
In wireless communication systems, parameter configuration has a profound impact on network performance. In particular, the rational setting of network parameters is crucial, as these parameters determine the design, energy consumption, and performance of the 6G network, and must be aligned with the real-time distribution of users and the electromagnetic environment. Based on the characteristics of wireless networks, network parameters can be categorized into two types: BS parameters and threshold parameters. Specifically, BS parameters involve basic switching functions such as VONR, paging, and slicing; threshold parameters configure the thresholds for cell reselection, event decision-making, and neighbor cell hand off. Moreover, spectrum resource planning in communication systems is also a critical component of parameter optimization. Studies have shown that spectrum is often underutilized in many time slots [80], and AI technologies have demonstrated great potential in cognitive radio [81]. AI methods can effectively monitor and mitigate interference [82], predict spectrum usage trends [83], and significantly improve the utilization of spectrum resources.
Conventional parameter optimization mainly relies on drive tests and expert experience. The current network parameter optimization methods are as follows: When network operators identify significant deficiencies in communication service quality, specific operators will deploy drive test vehicles to collect and report network performance metrics, such as network throughput and audio/video transmission delay. Subsequently, operators create scatter plots to further characterize network performance, and finally, expert experience is used to analyze and optimize the network parameters [84]. With the increasing complexity of network structures, conventional parameter optimization becomes inefficient. Due to the significant increase in the number of 6G antennas, the channel is no longer a simple one-dimensional circuit loss structure but has expanded into an N-dimensional angle power spectrum, resulting in a vast number of parameter combinations. The efficiency of expert-based parameter tuning is low, making it difficult to modify billions of 6G network parameters. Furthermore, conventional experience-based tuning methods often rely on inherent rules, which makes it difficult to fully capture and adapt to the complex relationships in the network [85]. Table 2 summarizes various parameters in the current networks.
@@
Currently, AI models have demonstrated unprecedented advantages in large-scale parameter tuning, nonlinear fitting, and real-time optimization [86]. Specifically, the increase in the number of antennas and the complexity of the 6G network has led to a surge in network parameter data, and AI can quickly search and optimize these combinations. In the context of the highly intricate architecture and nonlinear traits of the 6G network [87], AI technology can gain a more profound understanding of the network’s nonlinear features by learning from and dissecting extensive network data, thereby determining a more optimal parameter configuration. Moreover, in the 6G network, user requirements and network conditions change swiftly. AI technology can promptly optimize network parameters based on real-time data and scenario necessities, attaining enhanced performance.
In recent related research arenas, deep RL (DRL) has been deployed for the optimization of network parameters [88]. Specifically, each performance metric within the current communication network is considered as a distinct RL state, and the adjustment operations carried out on network parameters are designated as diverse actions. In this procedure, by incessantly adapting the behavior of the RL model across different states, the model will receive feedback information from the environment to assess the merit of each action. RL models typically focus on reinforcing those actions that can yield higher reward returns. Once the reward cost stabilizes, DRL will progressively approach a convergent state and subsequently determine the optimal configuration scheme of wireless network parameters.